Given that: depreciation = 90, indirect taxes = 70, subsidies = 30, payments to factors of production from abroad = 20, payments to foreign factors = 40.

Compute:

1) Gross national product (GNP) at market prices. (1 Mark)

2) Net national product (NNP) at market prices. (1 Mark)

3) Net national product (NNP) at factor cost. (1 Mark)

4) Net domestic product (NDP) at factor cost (1 Mark)
1
Expert's answer
2018-11-09T15:42:09-0500

Given that: depreciation D = 90, indirect taxes = 70, subsidies = 30, payments to factors of production from abroad = 20, payments to foreign factors = 40.

1) Gross national product GNP = GDP + payments to factors of production from abroad - payments to foreign factors = GDP + 20 - 40 = GDP - 20.

2) Net national product NNP = GNP - D = GDP - 20 - 90 = GDP - 110.

3) Net national product NNP at factor cost = GNP - indirect taxes - D = GDP - 20 - 70 - 90 = GDP - 180.

4) Net domestic product NDP at factor cost = GDP - indirect taxes - depreciation = GDP - 70 - 90 = GDP - 160.